I am confused by the discussion during the call today. I revisited Hartmut's paper - the circuit in Fig 6 is essentially what I am calling "only 2f_2 notch stuffed" in my previous elog. Qualitatively, the plot I presented in Attachment #2 of the preceeding elog in this thread shows the expected behavior as in Fig 8 of the paper - the impedance seen by the photodiode is indeed lower. In Attachment #1, I show the comparison - the "V(anode)/I(I1)" curve is analogous to the "PD anode" curve in Hartmut's paper, and the "V(vout)/I(I1)" curve is analogous to the "1f-out" curve. I also plot the sensitivity analysis (Attachment #2), by varying the photodiode junction capacitance between 100pF and 200 pF (both values inclusive) in 20 pF steps. There is some variation at 55 MHz, but it is unlikely that the capacitance will change so much during normal operation?
I understand the motivation behind stuffing the other notches, to reduce intermodulation effects. But the impression I got from the call was that somehow, the model I presented was wrong. Can someone help me identify the mistake?
I didn't bother to export the LTspice data and make a matplotlib plot for this quick analysis, so pardon the poor presentation. The colors run from green=100pF to grey=200pF.
On the call last week, I claimed that there isn't much hope of directly measuring Ponderomotive Squeezing in aLIGO without some significant configurational changes. Here, I attempt to quantify this statement a bit, and explicitly state what I mean by "significant configurational changes".
The I/O relations will generally look something like:
The. magnitudes of the matrix elements C_12 and C_21 (i.e. phase to amplitude and amplitude to phase coupling coefficients) will encode the strength of the Ponderomotive squeezing.
For the inital study, let's assume DC readout (since there isn't a homodyne readout yet even in Advanced LIGO). This amounts to setting in the I/O relations, where the former angle is the "homodyne phase" and the latter is the "SRC detuning". For DC readout, the LO quadrature is fixed relative to the signal - for example, in the usual RSE operation, . So the quadrature we will read out will be purely (or nearly so, for small detunings around RSE operation). The displacement noises will couple in via the matrix element. Attachment #1 and Attachment #2 show the off-diagonal elements of the "C" matrix for detunings of the SRC near RSE and SR operation respectively. You can see that the optomechanical coupling decays pretty rapidly above ~40 Hz.
In this particular case, there is no benefit to detuning the SRC, because we are assuming the homodyne angle is fixed, which is not an unreasonable assumption as the quadrature of the LO light is fixed relative to the signal in DC readout (not sure what the residual fluctuation in this quantity is). But presumably it is at the mrad level, so the pollution due to the orthogonal anti-squeezed quadrture can be ignored for a first pass I think. I also assume ~10 degrees of detuning is possible with the Finesse ~15 SRC, as the linewidth is ~12 degrees.
To see how this would look in an actual measurement, I took the data from Lee's ponderomotive squeezing paper, as an estimate for the classical noises, and plotted the quantum noise models for a few representative SRC detunings near RSE operation - see Attachment #3. The curves labelled for various phis are the quantum noise models for those SRC detunings, assuming DC readout. I fudged the power into the IFO to make my modelled quantum noise curve at RSE line up with the high frequency part of the "Measured DARM" curve. To measure Ponderomotive Squeezing unambiguously, we need the quantum noise curve to "dip" as is seen around 40 Hz for an SRC tuning of 80 degrees, and that to be the dominant noise source. Evidently, this is not the case.
The case for balanced homodyne readout:
I haven't analyzed it in detail yet - but it may be possible that if we can access other quadratures, we might benefit from rotating away from the DARM quadrature - the strength of the optomechanical coupling would decrease, as demonstrated in Attachments #1 and #2, but the coupling of classical noise would be reduced as well, so we may be able to win overall. I'll briefly investigate whether a robust measurement can be made at the site once the BHD is implemented.
Last week and this week I've been working on the characterization of the Q3000 QPDs. The QPDs were named 81, 82, 83, and 94.
My recommendation is to use #81 and #84 as they have similar dark current characteristics between the segments. But basically, all the QPDs look fine.
The actual junction capacitance and the RF dark noise should be characterized by the actual WFS head circuit.
The QPD packages were labeled and returned to Gautam to be implemented in the WFS heads.
gautam: S/N #84 was installed as the AS WFS QPD. The remaining 3 are stored in the clean cabinet at EX (where the rest of the RF photodiodes are).
Five Agilent pressure gauges were delivered to the 40m. It is stored with the controller and cables in the office area. This completes the inventory for the gauge replacement - we have all the ordered parts in hand (though. not necessarily all the adaptor flanges etc). I'll see if I can find some cabinet space in the VEA to store these, the clutter is getting out of hand again...
in addition, the spare gate valve from LHO was also delivered today to the 40m. It is stored at EX with the other spare valves.
It is stored along with the cables that arrived a few weeks ago, awaiting the gauges which are now expected next week sometime.
Basically, they repeated our specs and showed the coating performances for HR/AR for 10deg P and PR/AR for 45deg P. There is no RoC measurement by the vendor.
Nevertheless, their RoC (paper) specs should be compared with our request.
Optics --> Cabinet at south end (Attachment #1)
Scanned datasheets--> wiki. It would be good if someone can check the specs against what was ordered.
Attachment #1 - Proposed mods for 40m RF freqs.
Attachment #2 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
Attachment #3 - Modelled TFs for the case where all the notches are stuffed, and where only the 2f notch is stuffed.
Any other red flags anyone sees before I finish stuffing the board?
WFS head and housing. Need to finalize the RF transimpedance gain (i.e. the LC resonant part), and also decide which notches we want to stuff.
Where do we want to install the interface and readout electronics for the AS port WFS? Options are:
There isn't much difference in terms of cable length that will be required - I believe the AS WFS is going to go on the AP table even in the new optical layout and not on the ITMY in-air oplev table?
The project requires a large number of new electronics modules. Here is a short update and some questions I had:
Approximately half of the assembly of the various electronics is now complete. The basic electrical testing of the interface chassis and demod chassis are also done (i.e. they get power, the LEDs light up, and are stable for a few minutes). Detailed noise and TF characterization will have to be done.
Ordered 11/16 from CDW, on PO# S492940, the high voltage Tripp Lite SMART5000XFMRXL for TP-1. Should be arriving in about a week.
yes, both problems can be fixed. Usually we just order some spare lead-acid batteries from SRS (Steve may have some spare ones somewhere). The DC offset often comes from a busted FET input. I bought 50 of those at one point - they're obsolete. Its also possible to replace the input stage with any old FET pair.
I'll handle the one with the offset if you leave it on my desk.
Shruti picked it up @4pm.
true. also try to choose a cap with a goow high frequency response. In the Electronics Noise book by Ott there's some graph about this. I bet you good do a Bing search and also find something more modern. Basically we want to make sure that the self resonance is not happening at low frequencies. Might be tought to find one with a good HF response, a high voltage rating, and > 1uF.
Yes. The datasheet has a recommendation circuit with 10uF caps. Companies are careful to show reproducible, reliably functional circuit examples on datasheets. So, if the caps are there you should try to replicate the design.
Additional bypass capacitors? I use 0.1 uF, 700V DC ceramic capacitors as bypass capacitors close to the leads of the PA95, as is recommended in the datasheet. Can adding a 10uF capacitor in parallel provide better filtering? I'm not sure if one with compatible footprint and voltage rating is readily available, I'll look around.
I now think the excess noise in this circuit could be coming from the KEPCO switching power supply (in fact, the supplies are linear, and specd for a voltage ripple at the level of <0.002% of the output - this is pretty good I think, hard to find much better).
All component references are w.r.t. the schematic. For this test, I decided to stuff a fresh channel on the board, with new components, just to rule out some funky behavior of the channel I had already stuffed. I decoupled the HV amplifier stage and the Acromag DAC noise filtering stages by leaving R3 open. Then, I shorted the non-inverting input of the PA95 (i.e. TP3) to GND, with a jumper cable. Then I measured the noise at TP5, using the AC coupling pomona box (although in principle, there is no need for this as the DC voltage should be zero, but I opted to use it just in case). The characteristic bump in the spectra at ~100Hz-1kHz was still evident, see the bottom row of Attachment #1. The expected voltage noise in this configuration, according to my SPICE model, is ~10 nV/rtHz, see the analysis note.
As a second test, I decided to measure the voltage noise of the power supply - there isn't a convenient monitor point on the circuit to directly probe the +/- HV supply rails (I didn't want any exposed HV conductors on the PCB) - so I measured the voltage noise at the 3-pin connector supplying power to the 2U chassis (i.e. the circuit itself was disconnected for this measurement, I'm measuring the noise of the supply itself). The output is supposedly differential - so I used the SR785 input "Float" mode, and used the Pomona AC coupling box once again to block the large DC voltage and avoid damage to the SR785. The results are summarized in the top row of Attachment #1.
The shape of the spectra suggests to me that the power supply noise is polluting the output noise - Koji suggested measuring the coherence between the channels, I'll try and do this in a safe way but I'm hesitant to use hacky clips for the High Voltage. The PA95 datasheet says nothing about its PSRR, and seems like the Spice model doesn't include it either. It would seem that a PSRR of <60dB at 100 Hz would explain the excess noise seen in the output. Typically, for other Op-Amps, the PSRR falls off as 1/f. The CMRR (which is distinct from the PSRR) is spec'd at 98 dB at DC, and for other OpAmps, I've seen that the CMRR is typically higher than the PSRR. I'm trying to make a case here that it's not unreasonable if the PA95 has a PSRR <= 60dB @100 Hz.
So what are the possible coupling mechanisms and how can we mitigate it?
What do the analog electronics experts think? I may be completely off the rails and imagining things here.
Update 2130: I measured the coherence between the positive supply rail and the output, under the same conditions (i.e. HV stage isolated, input shorted to ground). See Attachment #2 - the coherence does mirror the "bump" seen in the output voltage noise - but the coherence is. only 0.1, even with 100 averages, suggesting the coupling is not directly linear - anyways, I think it's worth it to try adding some extra decoupling, I'm sourcing the HV 10uF capacitors now.
I had to go through five SR560s in the lab yesterday evening to find one that had the expected 4 nV/rtHz input noise and worked on battery power. To confirm that the batteries were charged, I left 4 of them plugged in overnight. Today, I confirmed that the little indicator light on the back is in "Maintain" and not "Charge". However, when I unplug the power cord, they immediately turn off.
One of the units has a large DC output offset voltage even when the input is terminated (though it is not present with the input itself set to "GND" rather than DC/AC). Do we want to send this in for repair? Can we replace the batteries ourselves?
The results from the ringdown are attached - in summary:
I want to test out an AS port WFS now that I have all the parts in hand - I guess the Michelson / PRMI will suffice until I make the ALS noise good again, and anyways, there is much assembly work to be done. Overnight, I'm repeating the suspension eigenmode measurement.
For the input matrix diagonalization, it seemed to me that when we had a significant seismic event or a re-alignment of the optic with the bias sliders, the input matrix also changes.
Meaning that our half-light voltage may not correspond to the half point inside the LED beam, but that rather we may be putting the magnet into a partially occluding state. It would be good to check this out by moving the bias to another setting and doing the ringdown there.
While proceeding with the interferometer recovery, I noticed that there appeared to be no light on WFS2. I confirmed on the AP table that the beam was indeed hitting the QPD, but the DC quadrants are all returning 0. Looking back, it appears that the failure happened on Monday 26 October at ~6pm local time. For now, I hand-aligned the IMC and centered the beams on the WFS1 and MC2T QPDs - MCT is ~15000 cts and MC REFL DC is ~0.1, all consistent with the best numbers I've been able to obtain in the past. I don't think the servo will work without 1 sensor without some retuning of the output matrix.
It would appear that both the DC and RF outputs of WFS2 are affected - I dithered the MC2 optic in pitch (with the WFS loop disabled) at 3.33 Hz, the transmission and WFS1 sensors see the dither but not WFS2. It could be that I'm just not well centerd on the PD, but by eye, I am, so it would appear that the problem is present in both the DC and RF signal paths. I am not going into the PD head debugging today.
Looking back through the elog, 1mtorr is the pressure at which it is deemed safe to send the full power beam into the IMC. After replacing the HR mirror in the MCREFL path with a 10% reflective BS, I just cranked the power back up. IMC is locked. With the increased exposure on the MC2T camera, lots of new scattered light has become visible.
I've uploaded some more photos here. I believe the problem is a worn out thread where the main rotary handle attaches to the shaft that operates the valve.
This morning, I changed the valve config such that TP2 backs TP1 and that combo continues to pump on the main volume through the partially open RV2. TP3 was reconfigured to pump the annuli - initially, I backed it with the AUX drypump but since the load has decreased now, I am turning the AUX drypump off. At some point, if we want to try it, we can try pumping the main volume via the RGA line using TP2/TP3 and see if that allows us to get to a lower pressure, but for now, I think this is a suitable configuration to continue the IFO work.
There was a suggestion at the meeting that the saturation of the main volume pressure at 1mtorr could be due to a leak - to test, I closed V1 for ~5 hours and saw the pressure increased by 1.5 mtorr, which is in line with our estimates from the past. So I think we can discount that possibility.
Main volume pressure as of 11:30AM 2020/11/10
[koji, rana, gautam]
1100 - EY chamber inspected, no issues were found --> EY heavy door on
1200 - OMC chamber was inspected. OM6 was marginally tweaked to bring the beam down a little in pitch, and also a little northwards in Yaw. --> Heavy door on.
1230 - Pumpdown started. Initially, the annuli volume was pumped down. The procedure calls for doing this with the small turbopumps. However, V7 was left open, and hence, in the process, the TP1 foreline pressure (=P2) hit ~30 torr. This caused TP1 to shutdown. We were able to restart it without issue. This case was not caught by the interlock code, which was running at the time. It should be recitified.
1330 - OMC breadboard clean optics and DCPD hardware were wrapped up and packed into tupperware boxes and stored along the south arm. OMC cavity itself, the OMMT, and the breadboard the OMC was sitting on are wrapped in foil/Ameristat and stored in cabinet S13, lower 2 shelves.
1915 - P1a = 0.5 torr pressure reached. Switched over to pumping the main volume with TP1, backed by TP2 and TP3, which themselves are backed by their respective dry pumps and also the AUX drypump for some extra oomph. All cooling fans available in the area were turned on and directed at the turbo pumps. RV2 was used to throttle the flow suitably.
It was at this point that we hit a snag - RV2 has gotten stuck in a partially open position, see Attachment #1. We can see that the thread doesn't move in response to turning the rotary dial. Fortunately, the valve is partially open, so the main volume continues to be pumped - see Attachment #2 for the full history of today's pumping. We are leaving the main volume pumped in this configuration overnight (TP1 pumping main volume backed by TPs 2 and 3, which are in turn backed by their respective drypumps and also the AUX dry pump). I think there is little to no risk of any damage to the turbo pumps, the interlocks should catch any anomalies. The roughing pumps RP1 and RP3 were turned off and that line was disconnected and capped.
What are our options?
We need some vacuum experts to comment. Why did this happen? Is this an acceptable failure mode of the valve?
2230 - P1a = 0.025 torr. The pressure is coming down with log-linear scale. x0.1 per 2.5 hours or so.
> I didn't bother to align the green beams to the arm cavities or re-center the Oplevs - is this necessary? It is a step in the pre-close up checklist, so maybe we should do it... The green transmission does reach the PSL table...
I don't think so. The beam is reaching the PSL, so we have no motivation to change the green alignment. Regarding the oplev, the green refl should come back to the PDH PD and this gives us additional beam reference. As soon as we find the green resonance after the pumping, we can tweak the green axis so that the spots on the mirrors become reasonable (as well as the green trans CCD on the PSL table).
Basic IFO alignment checks were done.
Tomorrow, we should do some visual checks of the chambers / EQ stops on ETMY etc but I don't see any major problems at the moment...
Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020.
I was able to boot one of the 3 new Supermicro machines, which I christened c1bhd, in a diskless way (with the boot image hosted on fb, as is the case for all the other realtime FEs in the lab). This is just a first test, but it is reassuring that we can get this custom linux kernel to boot on the new hardware. Some errors about dolphin drivers are thrown at startup but this is to be expected since the server isn't connected to the dolphin network yet. We have the Dolphin adaptor card in hand, but since we have to get another PCIe card (supposedly from LLO according to the BHD spreadsheet), I defer installing this in the server chassis until we have all the necessary hardware on hand.
I also have to figure out the correct BIOS settings for this to really run effectively as a FE (we have to disable all the "un-necessary" system level services) - these machines have BIOS v3.2 as opposed to the older vintages for which there are instructions from K.T. et al.
There may yet be issues with drivers, but this is all the testing that can be done without getting an expansion chassis. After the vent and recovering the IFO, I may try experimenting with the c1ioo chassis, but I'd much prefer if we can do the testing offline on a subnet that doesn't mess with the regular IFO operation (until we need to test the IPC).
I am working on the setup of a CDS FE, so please do not attempt any remote login to the IPMI interface of c1bhd until I'm done.
The PMC servo railed and so I re-locked it at ~half range. I've been noticing that the diurnal drift of the PZT control voltage has been larger than usual - not sure if it's entirely correlated with temperature on the PSL table. Anyway the cavity is locked again so all is good.
Attachment #1 shows the main result - there are 4 peaks. The frequencies are a little different from what I have on file for ETMY and the Qs are a factor of 3-4 lower (except SIDE) than what they are in vacuum, which is not unreasonable I hypothesize. The fits suggest that the peak shape isn't really Lorentzian, the true shape seems to have narrower tails than a Lorentzian, but around the actual peak, the fit is pretty good. More detailed diagnostic plots (e.g. coil-to-coil TFs) are in the compressed Attachment #2. The condition number of the matrix to diagonalize the sensing matrix (i.e. what we multiply the "naive" OSEM 2 Euler basis matrix by) is ~40, which is large, but I wouldn't read too much into it at this point.
I see no red flags here - the PIT peak is a little less prominent than the others, but looking back through the elog, this kind of variation in peak heights doesn't seem unreasonable to me. If anyone wants to look at the data, the suspension was kicked every ~1100seconds from 1288673974, 15 times.
I'm measuring the free-swinging spectra of ETMY overnight.
So all the primary vent objectives have been achieved 🙌 . The light doors are on the chamber right now. I'm measuring the free-swinging spectra of ETMY overnight. Barring any catastrophic failures and provided all required personnel are available, we will do the final pre-close-up checks, put the heavy doors back on, and pump down starting 10 am Monday, 9 Nov 2020. Some photos here.
The IMC isn't resonant for a TEM00 mode at the time of writing - we are waiting for the stack to relax, at which point if the IMC isn't resonant for a TEM00 mode, we will tweak the input pointing into the IMC (we want to use the suspended cavity as the reference, since it is presumably more reliable than the table from which we removed ~50 kgs of weight and shifted the balance.
I got a call from Calum ~830am today saying some facilities people entered the lab, opened the south entrance door, and tripped the alarm in the process. I came to the lab shortly after and was able to reset the alarm by flipping the switch on the alarm box at the south end entrance to "Alarm OFF". Then, I double checked that the door is closed, and re-enabled the alarm. The particle count at the SP table is not unusually high and the lasers (Oplev HeNe and AUX X) were still on, so doesn't look like any lasting damage was done. The facilities people were apparently wearing laser safety goggles.
Good point - looking back, I also see that I already removed the mirror at the SW corner of the table in 2016. Revised photo in Attachment #1. There is an optic on the east edge of this table whose purpose I'm not sure of, but I'm pretty sure it isn't essential to the main functionality and so can be removed.
I believe the mirror next to IM1 is for the green beams to be delivered to the PSL table. I think we still want to keep it. Otherwise, the plan looks fine.
To be a bit more clear about what we are going to do in the OMC chamber, I marked-up some photos, see Attachments #1 and #2.
I anticipate that after this work, the only components on the table will be
Are we in agreement with this plan?
See #15656 for the updated photo
This morning, we did the following;
The OSEMs remain in the EY vacuum chamber. The next set of steps are:
We will most likely work on this tomorrow. At ~1615, I briefly opened the PSL shutter and tweaked the IMC alignment. We will almost certainly change the pointing into the IMC when we remove the old OMC and rebalance that table, so care should be taken when working on that...
We are now ready to take the doors off. I've already done the basic prep work (loosened bolts, cleaned chamber, carts for tools, fresh ameristat on portable HEPAs etc).
If everything else looks good, I'll start letting the dry N2 into the main volume after lunch.
Now the green transmissions are visible by the green PDs. Attachment 9 shows the trans and ref of each green beams with and without locking to TEM00. The questionable green TRY was ~0.3. If we compare this with the histrical data (Attachment 10), it is about 1/4 of the value in the past. It's not too crazy but still quite low.
BTW, nice video! @ Koji, How difficult was it to edit it into this form?
I went to 40m yesterday at around 2:30 pm and Koji showed me how to acquire lock in different arms and for different lasers. Finally, we took a preliminary measurement of shaking the ETMX at some discrete frequencies and looking at the beatnote frequency spectrum of X-end laser's fiber-coupled IR and Main laser's IR pick-off.
We verified that we can send discrete frequency excitation signals to ETMX actuators directly and see a corresponding peak in the spectrum of beatnote frequency between fiber-coupled X-end IR laser and main laser IR pickoff.
If full interferometer had been locked, we could have used the DARM error signal output to calibrate it against this measurement.
At this point, I'm leaving the lab. All the suspensions (incl SRM) are aligned. PSL/GRX/GRY shutters were left open.
Ok I was using the PD in the black mount because Rana recommended it a few weeks ago.
Regarding the M2ISS, I acquired the hardware from QIL some months ago, including a circuit board, and 2 PDs. These had LEMO outputs though (not BNC), and the mounts are not 4". These photodiodes are what I'm using as the airBHD DCPDs right now, and some photos are here - are these the photodiodes you mentioned? Or are there yet more M2ISS photodiodes? I remember Johannes had some custom mounts extruded to make them 4" high, do you mean those? Can I retrieve them his Cryo setup?
BTW, my elog scraping shows only one spectra from Stefan in the ATF elog, and the performance there is more like 1e-7/rtHz @ 100 Hz, and that’s using a dedicated high BW servo circuit, not the SR560. Am I just missing the measurement of 2e-8/rtHz?
that little PD in the black mount was never very good. The AD829 is not a good opamp for transimpedance and especially not good for low frequencies. Stefan Ballmer and I were able to get 2e-8 out of these (@100 Hz) many years ago.
I wonder if we have some of Zach's M2ISS photodetectors around, perhaps in QIL or Cryo. I doubt that any of them are in use now. Those had good performance nad BNC output.
I'm starting the model restarts from remote. Then later I'll show up in the lab to do more hard resets.
==> It seems that the RFM errors are gone. Here are the steps.
I wanted to look into the ISS situation. Some weeks ago, I found the PD that was previously used as the in-loop photodiode. I wanted to use this and measure the open-loop RIN at a few places (to see if there's any variation and also to check its functionality). However, I didn't get very far tonight - for a start, the PD height is 3" (while our beam height is 4" everywhere outside the vacuum), and I needed to put together a circuit to supply the 5V bias and +/- 15 V since the transimpedance is done on the head. I was only able to do a low-level functionality test tonight, checked that the DC voltage output varied linearly with the incident power (calibrated against an NF1611 photodiode, data will be put up later). I didn't get to measuring any noise performance - is an incandescent light bulb still shot noise limited at ~10 Hz < f < 10kHz? Some notes:
Unconnected to this work - this problem reared its ugly head again (i noticed it yesterday morning already actually). I don't have the energy to embark on a fix tonight, Koji is going to be in the lab all day tomorrow and so he will fix it.
Apart from the questionable wiring on the Acromags, one other important difference is in the way the connections were made between the old VME crates to the Eurocrate backplanes, and how we do it now. The thick cables had their sheilds connected to the eurocrate ground (or at least, there was a dedicated ground lug on those cables which we screwed on to the ground terminals on the Eurocrate backplanes). However, in our current configuration, we interface the Acromag ADCs and DACs to the backplane via these adaptor boards. The shields of the DSUB cables are presumably NOT connected to the Eurocrate grounds. This should also be investigated as one potential cause of the grounding issue - while on some of the Eurocrate modules, the P1/P2 connectors may have either the "A" or "C" row of connectors shorted to ground, some may not, and the TTFSS may suffer from such an issue?
Note that we have this problem in all of the slow machines that were upgraded to Acromag (if this turns out to be the issue).
In fact, the problem was the grounding issue (presumably on the IOO racks).
In fact, the problem was the grounding issue (presumably on the IOO racks).
A temporary differential receiver at the TTFSS side was built using an SR560 and a few ponoma cables. This removed the structures ~850Hz.
The MC Servo Output was disconnected from the TTFSS box and monitored with SR785. The 850Hz structure was kept visible no matter what cables, including all the acromag DB cables, were removed. This made me suspicious about the measurement setup. The SR785 was connected to an AC power strip under the SP table and this was too far from the IOO rack.
The SR785 was connected to the AC power strip on 1X2, and now the difference becomes clear. No matter if the acromag cables are connected or not, the connection (particularly ground connection) between the MC servo module and the TTFSS box causes the MC servo output contaminated. (Comparison between Blue and Orange of Attachment #1). During the measurement, the EPICS switch for the fast path was disengaged (=no signal) and the VCO gain (...so called. It's just the MC Servo Gain) was set to be 0dB.
To test if the differential receiving of the MC Servo Output at the PSL helps to reduce this noise, I've built a simple (hacky) differential receiver using an SR560. (Attachment #2)
This kept the noise level same as the disconnected case (Comparison between Green and Orange of Attachment #1, I don't think the difference between them is not significant), while the IMC is locked as before.
Note that we can see that the 36kHz line was significantly reduced. Did we remove this annoying noise?
After talking with Gautam, we decided to leave this configuration while the SE-Diff cable was replaced with a more robust one. (See Attachment #3)
The PSL laser frequency performance was evakluated in the following two ways as we did last week:
1) Use the beat frequency of the free running PSL and the Y-end laser (Attachment #4). The PSL shutter was closed and thus the IMC was not locked.
2) Use the IMC MCF while the IMC was locked. (Attachment #5)
For both cases, the improvement was confirmed.
I also tried to check the reported issue by Gautam on this elog. He used 1Hz BW, but I cheated with 16Hz BW and 10x12.8kHz span PSDs. (Attachment #6)
For the measurement, IN1 GAIN of the IMC Servo was set to be 0dB and the OUT2 was switched to monitor the IN1 noise, while IN1 was terminated by a 50Ohm.
As I mentioned above, the AC power of SR785 was taken from a 1X2 power strip. Is this the reason for the power line forest look less severe compared to the previous case???
Anyway, I tried to use the same differential receiving technique (but with gain of x100) to see if this helps. The differential receiver helped to reduce the structure above 50kHz. The floor noise level was observed to be higher. I didn't pursue this any further, but the forest of the power line looked like a part of the measurement noise. This is indicative that the grounding condition on 1X2 is really not great and we need to review the configuration of the acromag grounding.
The particle counter on the 40m PSL was removed. The package was made together with the OMC lab particle counter (see the packing list below).
The kit was picked up by Radhika for a python code to read out the numbers.
=== Packing List ===
We wanted to track down the excess noise seen in MC_F and other places (see the previous report by Gautam)
Setup1: The IMC was locked and MC_F signal between 500 and 1500Hz was observed. The DTT template was saved as /users/Templates/MC/MCF_noise_201023.xml
- Suspected mech resonance/jitter coupled with clipping or any other imperfections. Poked the various optics and optomechanics on the table. Basically no change. If we tap the laser chassis and the optics close to the laser source, we occasionally unlocked the IMC
- When we touched (lifted) the Innolight controller box from the shelf, for the first time we saw a significant change in the shape of the noise spectrum. The peak around the 700Hz shited towards lower frequency by a few %. Other peaks have no obvious change in the shapes and the heights.
- While observing the MC_F signal on the laptop, we went to the back of the laser controller. Placing a hand close to the fan clearly changes the peak frequency lower. By temporarily disconnecting the fan from the power supply for a short moment, the 700Hz peak could be eliminated. We also tried to see the noise level with the slow thermal servo and diagnosis DB cable disconnected, but we didn't see any significant change of the noise level.
Setup 2: Using the ALS phase tracker, we can observe the relative freq noise of the PSL laser and the ETMY AUX laser without any servo involved. This way we can freely disconnect any cables from the lasers. The measurement template for DTT was saved as /users/Templates/ALS/Y_ALS_FINE_PHASE_OUT_102320.xml
- Noise spectrum before disconnecting the cable (REF0, RMS REF1)
- The Fast PZT input to the PSL was disconnected => This made all the peaks (including the 700Hz) disappeared (REF2, RMS REF3)
- The Fast PZT input was restored as before, then the chain was disconnected at the input of the HV PZT driver (Thorlabs) => Again, this made the peaks disappeared (REF4, RMS REF5)
- The chain was disconnected at the input of the TTFSS box => Again, this made the peaks disappeared (REF6, RMS REF7)
- Disconnected the demod input and the AO cables from the IMC servo board => This made the peaks came back (REF8)
- Disconnected all the input/peripheral cables from the IMC servo board except for the connection to the TTFSS box => Still the excess noise was observed (REF9)
- In addition to the above, the cable to the FSS box was disconnected but the ground was still touching the MC servo board => This made the peaks disappeared (REF10)
The conclusion is that the noise is injected from the main circuit of the IMC servo board.
Next time we will check if the backplane connection is doing something wrong. Also, we'll test if the presence of the RF signals does something bad to the IMC board via EMI and RFI.
We have reverted the connection and tested if we lock the IMC and Y arm. ==> We saw at least they were locked for a short period. The things are still stabilizing, but left them turned on so they keep trying to lock automatically for the night.
Andrew made a battery-powered 0.7 nVrtHz input-referred noise pre-amplifier for gain of 200. That might help you.
we'd need a preamp with better than 1nV/rtHz to directly measure the noise I guess.
RXA: 0.7 nV is OK if you're not interested in low noise measurements. Otherwise, we have the transformer coupled pre-amp from SRS which does 0.15 nV/rHz and the Rai Weiss FET amp which has 0.35 nV for high impedance sources.